Find the mid-band gain of the circuit AM as well as the corner frequencics due to parasitic capacitance fz, f., and fir. Assume f-150), VBE 0.7 V, VA = oo V, Cr = 35 pF, and C,, = 4 pF 10 V 5 k2 42.5 k23 C2 10 kS2 02 20.5 kg 0.1 kΩ CCI 21 28.3 kΩ 5.4 kS2 -10 V The necessary formulas are given below: 1- A BJT Emitter Bias BJT Relationships and V C,max VC.mx t VC min SS Hybrid-T BJT Model R10 X RE al enors CTORY ual errors PING strabed, but SFACTORY critically p esent the is nada
Answer #1

VP RC C3 R1 Vout RL - C1 Q2 R2 C2 Vin -w Q1 RS R3 RE1 CE1 0.1 uF

R1

42.5 (K ohms)

R2

20.5(K ohms)

R3

28.3(K ohms)

RC (Collector resistor)

5(K ohms)

RE1 (Emitter resistor, AC bypassed)

0(K ohms)

RE2 (Emitter resistor, gain control)

5400(ohms)

VP (Supply Voltage)

10(V)

Beta (DC Current Gain)

150

VBE (Base to emitter drop)

0.7(V)

Rs (Source Resistance):

100(ohms)

RL (Load resistor)

10000(ohms)

CCB (Cu Collector-Base Cap.):

4 (pF)

CBE (Cπ, Base-Emitter Cap.)

35(pF)

VB1= R3/(R1+R2+R3) =4.645 V

VB2= (R2+R3)/(R1+R2+R3) = 3.0997 V

VE1= VB1-VBE = 2.3997 V

VE2= VB2-VBE = 4.6450 V

VC1= VE2=4.6450 V

VC2= VP - IC*RC=7.8304 V

RB1= R3*(R1+R2)/ (R1+R2+R3)

IB*RB1 + IB*Beta*RE+VBE = VB1

IB= (VB1-VBE)/(RB1+Beta*RE)

IC= Beta*IB =0.43392 mA

gm = Ic/25mA =0.017357

rπ= Beta/gm = 8642.1 Ohms

If RE2 is zero ohms:

A= -gm*((RL||RC)*(R2||R3))/(R23+RS) * ( rπ/( rπ+Rx+(R2||R3||RS)),

else

A= -(RL||RC)/RE2 =-0.61728

Rx is typically10-30 ohms, we use 20 in the calculations.

fH= gm/(2π*(CBE+CBC)) = 70.86 MHz

f1= 1/(Rin*(Cbe+ 2*Ccb))/(2*π)= 157.37 MHz

f2= 1/(RL*Ccb)/(2*π) = 11.937 MHz